OpenStudy (anonymous):

what is 0 times 1/0?

OpenStudy (juanpablojr):

0 times 1 is 0 / 0 = 1 ?

OpenStudy (acxbox22):

1/0 is undefined so if you multiply it by 0 it will remain undefined

OpenStudy (uri):

syntax error

OpenStudy (xapproachesinfinity):

Error!

OpenStudy (xapproachesinfinity):

lol

OpenStudy (solomonzelman):

1/0 is undefined for sure, and 0/0 is on some level equivalent to 1. So when you say `0 × (1/0)` you are multiplying the number that is for sure undefined by 0.

OpenStudy (solomonzelman):

0/0 can be 1 or undefined 0/1 can ONLY be UNDEFINED

OpenStudy (solomonzelman):

For your question, just put "undefined"

OpenStudy (anonymous):

ok thank you but so your saying it can also be 1?

OpenStudy (xapproachesinfinity):

haha

OpenStudy (solomonzelman):

Well, on some level. It is very complicated. You can sa that 0/0 is any number technically, because when you look at `0/0 = d/1` and cross multiply, you get ` 0 × d = 0 × 1` which gives `0=0`. See my point? Well, it is fair to say though, that UNDEFINED is the BEST answer.

OpenStudy (solomonzelman):

So when you prove from 0/0=d/1 you can say that d is any number, in other words it is undefined, just because it can be any number.

OpenStudy (solomonzelman):

it can be ∞

OpenStudy (anonymous):

ok thanks so much for explaining it so well. I was asking this question because i was looking a somones anser on open study for my real question and if you want to help me a little more if you can take a look at this link. scroll down to scrollpatrols anser : http://openstudy.com/study#/updates/52de7fbce4b003c643a11055

OpenStudy (anonymous):

it involves the question 0 times !/0

OpenStudy (solomonzelman):

It can't be ( 0 × ! ) / 0 it has to be a factorial of something, or 0! / 0 perhaps ?

OpenStudy (solomonzelman):

0! / 0 is same as 1/0, because 0! is defined to be 1.

OpenStudy (anonymous):

oh ok so do you think that he used a good example on explaining the question?

OpenStudy (solomonzelman):

Well, saying that 0 × (1/0) = 1 if you cancel the top and above zero, but there is a rule that `anything times zero, is equivalent to zero` I can sort of agree with canceling the denominator with the number you multiply by (which in this case is a zero), but I really do not relate multiplying anything times 0/1? Because 0/1 × 0 is same as 'undefined' × 0, and shouldn't be equal to zero or 1. I would say use the one I sad about 0/0 = d/1 (because 0/0 is at least somehow a number) But try it for 0/4. If you have 0 pizza slices, and 4 friends, and you distribute those 0 pizza slices (you're feeling generous...) equally amongst each of them, how much would each of your 4 friends get? Clearly, they would each get 0 dollars! Now try it for 4/0. If you have 4 pizza slices and 0 friends, and you.... but how can you distribute any amount of money amongst friends who don't exist? So the question of what 4/0 means makes no sense!

OpenStudy (solomonzelman):

bottom line, 0/0 and 0/1 are both undefined.

OpenStudy (solomonzelman):

Another way: 0/0 (like distributing 0 piece of cakes amongst 0 friends), so you can have this equivalent to any number, because you just not distributing anything to anyone. so, 0/0=15 check it, 15 × 0 =0 Nope, check doesn't work. Extraneous solution :)

OpenStudy (solomonzelman):

same way 0/0 can be equal to any negative or positive real number. And perhaps even to an imaginary number.

OpenStudy (anonymous):

ok thank you so much. now i just have to figure out which example i want to use to slove my original question about kelly

OpenStudy (solomonzelman):

Original question about Kelly ?

OpenStudy (anonymous):

its the question that i posted the link to. That my original question that i need to slove

OpenStudy (solomonzelman):

`Kelly tells you that when variables are in the denominator, the equation becomes ` `unsolvable. "There is a value for x that makes the denominator zero, and you can't` ` divide by zero," Kelly explains. Using complete sentences, demonstrate to Kelly ` `how the equation is still solvable.` There is a value of a variable that makes a denominator equal zero, and this value is called a restriction. IF this variable is x, then it is a domain restriction, and if it y or f(x) then it is a range restriction. You don't want to have 0/0 because it is undefined. She is wrong however in her first line, because if there is a variable in a denominator, then there will be a restriction, but there will be a solution. You can solve the equation (last sentence), however it depends on the equation if it will have x (or y) intercept(s) .

OpenStudy (anonymous):

oh ok im kinda getting it. so when it is asking me that it is solvable i would say that it depends on the equation if it will have x(ory) intercepts?

OpenStudy (solomonzelman):

Yes, you would say that it is typically solvable. (by the 'solutions' they are probably referring to x-intercept(s) )

OpenStudy (anonymous):

so they are saying to demonstrate to kelly that it is solvable

OpenStudy (solomonzelman):

They are asking you to explain it in general i.e. without using any specific example. (Which is harder)

OpenStudy (solomonzelman):

You should propose some explanation/theory to this.

OpenStudy (solomonzelman):

Well, if Kelly had a different equations that SHE was asking about, then you would show it on this equation.

OpenStudy (anonymous):

yeah your right. ok i think ill put something together to anser the question. Thanks so much for your time and patients and your help. (:

OpenStudy (solomonzelman):

Anytime ! enjoy ;)